Hello Tutors,
Can you please help me to solve the following problem?
A card is drawn from a​ well-shuffled deck of 52 cards. Find the probability that the card is a black card or a 5.
Thank you in advance

Answer by rothauserc(4718)

there are 26 black cards in the deck and 4 5's
:
probability of drawing a black card is 26/52 = 1/2
:
probability of drawing a five is (4-2)/52 = 2/52 = 1/26
:
Note we have to subtract the 2 black fives since they are already counted in the probability for a black card
:
probability of drawing a black card or a five is (1/2) + (1/26) = 14/26 = 7/13
: